OpenStudy (anonymous):

3. A shipping company designs a container that is a rectangular prism with a length that is 3 inches more than its width and a height that is 1 inch less than its width. If the volume of the container is 160 cubic inches, find its dimensions. Show all of your work.

OpenStudy (faiqraees):

lets assume width = x and thus length will be = x+3 and height = x-1

OpenStudy (faiqraees):

Any confusion

OpenStudy (anonymous):

okay I understand so far

OpenStudy (anonymous):

so what am I supposed to do an equation or sum?